# latexpage numbers missing

### v1

@@ 5 Acknowledgments
$\bm{x}_{t-1}\approx\big{(}1+\frac{\beta_{t}}{2}\big{)}\big{(}\bm{x}_{t}+\beta_{t}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)\big{)}+\sqrt{\beta_{t}}\bm{\epsilon}$
$\bm{x}_{t-1}\approx\big{(}1+\frac{\beta_{t}}{2}\big{)}\bm{x}_{t}+\beta_{t}\big{(}1+\frac{\beta_{t}}{2}\big{)}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)+\sqrt{\beta_{t}}\bm{\epsilon}$
$\bm{x}_{t-1}\approx\big{(}1+\frac{\beta_{t}}{2}\big{)}\bm{x}_{t}+\beta_{t}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)+\frac{\beta_{t}^{2}}{2}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)+\sqrt{\beta_{t}}\bm{\epsilon}$
$\bm{x}_{t-1}\approx\big{(}1+\frac{\beta_{t}}{2}\big{)}\bm{x}_{t}+\beta_{t}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)+\sqrt{\beta_{t}}\bm{\epsilon},\quad\textrm{since }\beta_{t}^{2}\approx 0$
$\bm{x}_{t-1}-\bm{x}_{t}\approx\frac{\beta_{t}}{2}\bm{x}_{t}+\beta_{t}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)+\sqrt{\beta_{t}}\bm{\epsilon}$
$\bm{x}_{t-1}-\bm{x}_{t}\approx-\bm{f}(\bm{x}_{t},t)+g(t)^{2}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)+g(t)\bm{\epsilon},\quad\bm{f}(\bm{x}_{t},t)=\frac{-\beta_{t}}{2}\bm{x}_{t},\quad g(t)=\sqrt{\beta_{t}}$
$\bm{x}_{t}-\bm{x}_{t-1}\approx\bm{f}(\bm{x}_{t},t)-g(t)^{2}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)+g(t)\bm{\epsilon},$
where the second line follows due to a MacLaurin series with $h(\beta_{t})=(1-\beta_{t})^{-\frac{1}{2}}$ for $\beta_{t}$ selected such that $\beta_{t}\approx 0$ for $t\in[0,T]$.
@@
We first applied NCS to a Gaussian process as a sanity check because conditional simulation from a Gaussian process is tractable and consequently, we can validate NCS using metrics which utilize the true predictive distributions.
@@
Here, we elaborate on the brief description of Brown–Resnick processes in Section 3.1. Max-stable processes can be expressed as the pointwise maximum of an infinite collection of random spatial processes (Davison et al., 2012; Huser et al., 2019). More specifically, a max-stable process has the following construction (de Haan, 1985)
$Z(\bm{s})=\max_{i>0}\eta_{i}W_{i}(\bm{s}),\textrm{ for }\bm{s}\textrm{ on a domain }\mathcal{D}\subset\mathbb{R}^{d},$ (S7)
in which $\{\eta_{i}\}$ is a Poisson point process on $(0,\infty)$ with intensity function $\operatorname{d}\!{\Lambda}(\eta)=\eta^{-2}\operatorname{d}\!{\eta}$ and $W_{i}(\bm{s})$ are independent copies of a non-negative stochastic process with mean one (Kabluchko et al., 2009). Different types of max-stable processes arise from the selection of the non-negative process $W_{i}(\bm{s})$. For a Brown–Resnick, the non-negative process $W_{i}(\bm{s})$ is a log-normal process of the form:
$W_{i}(\bm{s})=\exp(\epsilon_{i}(\bm{s})-\gamma(\bm{s})),$ (S8)
where $\epsilon_{i}(\bm{s})$ are realizations of an intrinsically stationary Gaussian process, conventionally with $\epsilon_{i}(\bm{s})=0$ almost surely, and with semivariogram $\gamma(\bm{h})$. A common choice, which we also adopt in this work is to set $\gamma(\bm{h})=(\left.\left\lVert\bm{h}\right\rVert\middle/\lambda\right.)^{\nu}$, with $\bm{h}$ the spatial separation between two locations, $\lambda\in\mathbb{R}^{+}$ a range parameter, and $\nu\in(0,2]$ a smoothness parameter (Castruccio et al., 2016).
S2.2 Training Details Common to Both Case Studies
Loss Function
We utilize the same weighting function as in Song et al. (2021), which is proportional to the inverse of the expectation of the conditional score function. This is equivalent to setting $\lambda(t)=1-\alpha_{t}$ for a variance preserving SDE (VPSDE):
$\lambda(t)$ $=\Big{(}\mathbb{E}\big{(}\left\lVert\nabla_{\tilde{\bm{x}}_{t}}\log(p(\tilde{\bm{x}}_{t}\mid\bm{x}_{0})\right\rVert_{2}^{2}\big{)}\Big{)}^{-1}$ (S9)
@@
For the parameter U-Net, we consider the parameter space $\Theta=[1,5]\times\{1.5\}$ in which $\nu=1.5$ is fixed for both processes and length scale $\ell$ and range $\lambda$ vary from one to five for the Gaussian process and Brown–Resnick process, respectively. The training parameter space $\tilde{\Theta}$—the parameter space from which we draw unconditional simulations during training—extends beyond $\Theta=[1,5]\times\{1.5\}$ to avoid suboptimal behavior at the boundary during evaluation. The proportion U-Net has the same parameter space for each spatial process: $\Theta=\{3\}\times\{1.5\}$ in which we fix the length scale $\ell=3$ and variance $\nu=1.5$ for the Gaussian process and the range $\lambda=3$ and smoothness $\nu=1.5$ for the Brown–Resnick process. Since the parameters are fixed for the proportion U-Net, the training parameter space $\tilde{\Theta}$ is the same as $\Theta$.
@@
For each data draw, we simulate a new dataset and use it to train and evaluate the diffusion model for a fixed number of epochs. This approach avoids overfitting and minimizes data storage (Chan et al., 2018).
We simulated training and validation data for the parameter and proportion U-Nets according to the following: For each data draw, we first simulate $p$ parameters from the training parameter space $\tilde{\Theta}$ via a uniform distribution and for each of those $p$ parameters, we simulate $s$ full spatial field realizations via the unconditional simulator. Finally, we simulate $m$ masks and $m$ discrete timesteps uniformly from $\{0,1,\dots,T\}\subset\mathbb{Z}$ per each full spatial field realization. Each mask is simulated using the Bernoulli distribution with $\rho$—the probability of observing a single location—for all $n=32^{2}$ locations. For the parameter U-Net, $\rho=0.05$ is kept fixed. For the proportion U-Net, $\rho$ is sampled uniformly from $[0.01,0.525]$ during training to avoid boundary issues during evaluation where the evaluation proportion space is $[0.01,0.5]$. Algorithm S1 illustrates this data generation process. Once finished with simulation, the data have the form:
@@
Algorithm S1 Simulating Training and Validation Data for Parameter and Proportion U-Nets
$\mathcal{D}=\emptyset$
Sample $q_{1},\dots,q_{r}\sim U([q_{L},q_{U}])$.
@@
for $\bm{\theta}_{j}$ in $\{\bm{\theta}_{j}\}_{j\in[l]}$ do
  $\{\bm{x}_{0,k}\}_{k\in[s]}\sim p(\cdot\mid\bm{\theta}_{i})$.
@@
   $\bm{M}(\mathring{\mathcal{S}}_{l})\sim B(n=32^{2},p=q_{i})$ for $l\in[m]$
   $t_{l}\sim U([0,T])\textrm{ for }l\in[m],\quad[0,T]\in\mathbb{Z}$
   $\bm{\mathring{x}}_{0,l}=\bm{M}(\mathring{\mathcal{S}}_{l})\odot\bm{x}_{0,k}\textrm{ for }l\in[m]$
   $\tilde{\bm{x}}_{t_{l}}\sim p(\cdot\mid\bm{\mathring{x}}_{0,l},\tilde{\bm{x}}_{0,k}):=\mathcal{N}(\sqrt{\tilde{\alpha}_{t}}\tilde{\bm{x}}_{0,k},\tilde{\sigma}_{t}^{2}\bm{I})\textrm{ for }l\in[m]$
   $\mathcal{D}\leftarrow\mathcal{D}\cup\{\big{(}(\bm{\mathring{x}}_{0,l},\tilde{\bm{x}}_{t_{l}}),\bm{M}(\mathring{\mathcal{S}}_{l}),\bm{\theta}_{j},t_{l}\big{)}\mid l\in[m]\}$
@@
We train U-Nets for the small conditioning sets in Section 3.2 via the following data generation process: We first fix the parameter $\bm{\theta}=(\lambda,\nu)$ and select a minimum $o_{L}$ and maximum $o_{U}$ number of observations for the conditioning sets. For each data draw, we iterate through each number of observed locations $o_{j}$ starting from the minimum $o_{L}$ and ending with the maximum $o_{U}$. For each observed number $o_{j}$, we simulate $s$ full spatial field realizations via the unconditional simulator with fixed parameter $\bm{\theta}$. For each of these $s$ full spatial fields, we construct $m$ masks and simulate $m$ discrete timesteps uniformly from $\{0,1,\dots,T\}\subset\mathbb{Z}$. Each mask is constructed from a conditioning set $\mathring{\mathcal{S}}$ containing $o_{j}$ locations uniformly sampled without replacement from $\mathcal{S}$. Algorithm S2 illustrates this data process, and the resulting data have the form shown in (S10).
Algorithm S2 Simulating Training and Validation Data for Small Conditioning Set U-Net

### v2

@@ Acknowledgments
$\approx\big(1+{\beta_{t}/2}\big)\big(\bm{x}_{t}+\beta_{t}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)\big)+\sqrt{\beta_{t}}\bm{\epsilon}$
$=\big(1+{\beta_{t}/2}\big)\bm{x}_{t}+\beta_{t}\big(1+{\beta_{t}/2}\big)\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)+\sqrt{\beta_{t}}\bm{\epsilon}$
$=\big(1+{\beta_{t}/2}\big)\bm{x}_{t}+\beta_{t}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)+\frac{\beta_{t}^{2}}{2}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)+\sqrt{\beta_{t}}\bm{\epsilon}$
$\approx\big(1+{\beta_{t}/2}\big)\bm{x}_{t}+\beta_{t}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)+\sqrt{\beta_{t}}\bm{\epsilon},\quad\textrm{since }\beta_{t}^{2}\approx 0,$
and therefore,
$\bm{x}_{t-1}-\bm{x}_{t}$ $\approx{\beta_{t}\bm{x}_{t}}/2+\beta_{t}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)+\sqrt{\beta_{t}}\bm{\epsilon}$ (S8)
$\approx-\bm{f}(\bm{x}_{t},t)+g(t)^{2}\bm{s}_{\bm{\phi}}(\bm{x}_{t},t)+g(t)\bm{\epsilon},$
where $\bm{f}(\bm{x}_{t},t)=-{\beta_{t}\bm{x}_{t}}/2$, $g(t)=\sqrt{\beta_{t}}$, and the first approximation follows due to a MacLaurin series expansion of $h(\beta_{t})=(1-\beta_{t})^{-\frac{1}{2}}$ for $\beta_{t}$ chosen such that $\beta_{t}\approx 0$ for $t\in[0,T]$.
@@
A Gaussian process is a stochastic process whose finite-dimensional distributions are Gaussian. Its predictive distributions are analytically available, and therefore we can use them to validate NCS. Results from applying NCS to Gaussian processes are shown in Section S3.3.
@@
Here, we give more details on the Brown–Resnick process employed in Section 3.1. Max-stable processes, such as Brown–Resnick processes, can be expressed as the pointwise maxima of an infinite collection of random spatial processes (Davison; Huser2019; Davison2019). More specifically, a max-stable process has the following construction (deHaan1985)
$Z(\bm{s})=\max_{i=1,2,\ldots}\eta_{i}W_{i}(\bm{s}),\textrm{ for }\bm{s}\textrm{ on a domain }\mathcal{D}\subset\mathbb{R}^{d},$ (S9)
in which $\{\eta_{i}\}_{i=1}^{\infty}$ is a Poisson point process on $(0,\infty)$ with intensity function $\operatorname{d}\!{\Lambda}(\eta)=\eta^{-2}\operatorname{d}\!{\eta}$ and $\{W_{i}(\bm{s})\}$ are independent copies of a non-negative stochastic process with mean one (Kabluchko2009). Different types of max-stable processes arise from the definition of $W_{i}(\bm{s})$. For a Brown–Resnick model, the process $W_{i}(\bm{s})$ is log-normal and has the form:
$W_{i}(\bm{s})=\exp(\epsilon_{i}(\bm{s})-\gamma(\bm{s})),\quad i\in\mathbb{N}^{+}$ (S10)
where $\{\epsilon_{i}(\bm{s})\}_{i\in\mathbb{N}^{+}}$ are realizations of an intrinsically stationary Gaussian process, conventionally with $\epsilon_{i}(\bm{s})=0$ almost surely, and with semivariogram $\gamma(\bm{h})$, $i\in\mathbb{N}^{+}$, where $\mathbb{N}^{+}$ denotes the positive integers. A common choice, which we also adopt in this work, is to set $\gamma(\bm{h})=(\left.\left\lVert\bm{h}\right\rVert\middle/\lambda\right.)^{\nu}$, with $\bm{h}$ the spatial separation between two locations in spatial domain $\mathcal{D}$ , $\lambda\in\mathbb{R}^{+}$ a range parameter, and $\nu\in(0,2]$ a smoothness parameter (Castruccio).
S2.2 Training Details Common to the Gaussian Process and Brown–Resnick Process Case Studies
Weighting Function
In our loss function (7a), we use the same weighting function $\lambda(\cdot)$ as in Song2021, which is proportional to the inverse of the expectation of the conditional score function. This is equivalent to setting $\lambda(t)=\underline{\sigma}_{t}^{2}$ for a variance preserving SDE (VPSDE), since, from (10),
$\lambda(t)$ $=\Big(\mathbb{E}\big(\left\lVert\nabla_{\tilde{\bm{x}}_{t}}\log(p(\tilde{\bm{x}}_{t}\mid\bm{x}_{0}))\right\rVert_{2}^{2}\big)\Big)^{-1}$ (S11)
@@
In the main text, we consider two differently amortized U-Nets, one amortized with respect to a spatial process parameter referred to as a parameter U-Net and the other amortized with respect to the proportion of missing observations referred to as a proportion U-Net. For the parameter U-Net, we consider the parameter space $\Theta=[1,5]\times\{1.5\}$, where $\nu=1.5$ is fixed for both processes described in Section S2.1, and where the length scale $\ell$ and range $\lambda$ vary from one to five for the Gaussian process and Brown–Resnick process, respectively. We define the training parameter space $\tilde{\Theta}$—the space from which we draw unconditional simulations during training—to extend beyond $\Theta=[1,5]\times\{1.5\}$ to avoid suboptimal behavior of the U-Net at the boundary. On the other hand, when training the proportion U-Net, we do not vary the spatial process parameters. Specifically, we fix the length scale $\ell=3$ and variance $\tau^{2}=1.5$ for the Gaussian process, and the range $\lambda=3$ and smoothness $\nu=1.5$ for the Brown–Resnick process. Therefore, for the parameter U-Net, $\Theta=\tilde{\Theta}=\{3\}\times\{1.5\}$.
@@
We employ data generation “on the fly.” That is, we simulate a dataset, use it to train and evaluate the diffusion model for a fixed number of epochs, and repeat the process with a new simulated dataset each time. This approach avoids overfitting and minimizes data storage (Chan2018).
We simulated training and validation data for the parameter and proportion U-Nets according to the following procedure: For each data draw, we first simulate $p$ parameters from the training parameter space $\tilde{\Theta}$, and for each of those $p$ parameters, we simulate $s$ full spatial field realizations via the unconditional simulator. Note that for the proportion U-Net, these $p$ parameter vectors are identical. Finally, we simulate $m$ masks and $m$ discrete timesteps uniformly from $\{0,1,\dots,T\}\subset\mathbb{Z}$ per each full spatial field realization. Each mask is simulated using the Bernoulli distribution with $\rho$—the probability of observing a single location—for all $n=32^{2}$ locations, and our convention is that 0 indicates missingness. For the parameter U-Net, $\rho=0.05$ is kept fixed. For the proportion U-Net, we sample $\rho$ uniformly from $[0.01,0.525]$ during training to avoid boundary issues during evaluation, where we instead restrict the space of proportions to $[0.01,0.5]$. Algorithm S1 illustrates this data generation process. Once finished with simulation, we have training data of the form:
$\{(\bm{\mathring{x}}_{0}^{\prime},\tilde{\bm{x}}_{t}^{\prime})^{\prime},\bm{M}(\mathring{\mathcal{S}}),\bm{\theta},t\}$ (S12)
@@
Algorithm S1 Simulating Training and Validation Data for the Parameter and Proportion U-Nets
Set $\mathcal{D}=\emptyset$
Parameter U-Net: Set $q_{L}=q_{U}=0.05$, $\tilde{\Theta}=[.5,5.5]\times\{1.5\}$, $\Theta=[1,5]\times\{1.5\}$.
Proportion U-Net: Set $q_{L}=0.01$, $\Theta=\{3\}\times\{1.5\}$, and $q_{U}=0.525$ for the training data and $q_{U}=0.5$ for the validation data.
Sample proportions of observed locations $q_{1},\dots,q_{r}\sim U([q_{L},q_{U}])$.
@@
for $\bm{\theta}_{j}$ in $\{\bm{\theta}_{j}\}_{j\in[p]}$ do
  Simulate unconditionally $\{\bm{x}_{0,k}\}_{k\in[s]}\sim p(\cdot\mid\bm{\theta}_{j})$.
@@
   $\bm{M}(\mathring{\mathcal{S}}_{l})\sim\textrm{Bin}(n=32^{2},p=q_{i})$ for $l\in[m]$
   $t_{l}\sim U(\{0,1,\dots,T\})\textrm{ for }l\in[m]$
   $\bm{\mathring{x}}_{0,l}=(x_{0,k,j}:M_{j}(\mathring{\mathcal{S}}_{l})=1)^{\prime}\textrm{ for }l\in[m]$
   $\tilde{\bm{x}}_{0,l}=(x_{0,k,j}:M_{j}(\mathring{\mathcal{S}}_{l})=0)^{\prime}\textrm{ for }l\in[m]$
   $\tilde{\bm{x}}_{t_{l}}\sim p(\cdot\mid\bm{\mathring{x}}_{0,l},\tilde{\bm{x}}_{0,l}):=\mathcal{N}(\sqrt{\underline{\alpha}_{t}}\tilde{\bm{x}}_{0,l},\underline{\sigma}_{t}^{2}\tilde{\bm{I}})\textrm{ for }l\in[m]$
   $\mathcal{D}\leftarrow\mathcal{D}\cup\{\{(\bm{\mathring{x}}_{0,l}^{\prime},\tilde{\bm{x}}_{t_{l}}^{\prime})^{\prime},\bm{M}(\mathring{\mathcal{S}}_{l}),\bm{\theta}_{j},t_{l}\big\}:l\in[m]\}$
@@
S2.3 Validation Details Common to Both the Gaussian Process and Brown–Resnick Process Case Studies
S2.3.1 Generation of Validation Datasets
In this section, we give details on how we generate test data for the simulation experiments conducted in Section 3 and Supplementary Materials S3 and S4.
Conditional Validation Dataset (Proportion U-Net)
For each of the five observed proportions $\rho\in\{0.01,0.05,0.1,0.25,0.5\}$, we generated a conditioning set $\mathring{\mathcal{S}}$ and mask $\bm{M}(\mathring{\mathcal{S}})$ using the Bernoulli distribution with probability $\rho$, and a reference spatial field $\bm{x}_{0}$ from its unconditional distribution with either variance $\tau^{2}=1.5$ and length scale $\ell=3$ (Gaussian process), or smoothness $\nu=1.5$ and range $\lambda=3$ (Brown–Resnick process). For each of the resulting five partially observed spatial fields $\bm{\mathring{x}}_{0}=(x_{0,j}:M_{j}(\mathring{\mathcal{S}})=1)^{\prime}$, we generated $m=4000$ replicates of the unobserved spatial field $\{\tilde{\bm{x}}_{0,i}\}_{i\in[m]}$ with NCS and either the true predictive distribution (Gaussian process) or an MCMC approximation (Brown–Resnick process). We use this dataset in Section 3.3 to validate NCS when the percentage of observed locations varies.
Conditional Validation Dataset (Parameter U-Net)
In the Gaussian process case, we unconditionally simulated a reference spatial field $\bm{x}_{0}$ for each parameter length scale $\ell\in\{1,\ldots,5\}$ and variance $\tau^{2}=1.5$. In the Brown–Resnick case, we unconditionally simulated a reference spatial field $\bm{x}_{0}$ for each range parameter $\lambda\in\{1,\ldots,5\}$ and smoothness $\nu=1.5$. Then, for each reference spatial field, we generated a conditioning set $\mathring{\mathcal{S}}$ and corresponding mask $\bm{M}(\mathring{\mathcal{S}})$ using the Bernoulli distribution with probability $\rho=0.05$ of observing each location. In each case, using each of the five partially observed fields $\bm{\mathring{x}}_{0}=(x_{0,j}:M_{j}(\mathring{\mathcal{S}})=1)^{\prime}$, we generated $m=4000$ unobserved spatial field replicates $\{\tilde{\bm{x}}_{0,i}\}_{i\in[m]}$ with NCS and either the true predictive distribution (Gaussian process) or an MCMC approximation (Brown–Resnick process). We use this dataset in Section 3.3 to validate NCS when the spatial process parameter varies.
S2.3.2 Diagnostic and Validation Metrics
In this section, we describe diagnostics and metrics used to assess the quality of conditional simulation in Sections 3.2, 3.3, S3.2 and S4.4.
Conditional Univariate and Bivariate Densities
Conditional univariate and bivariate densities are estimated using the conditional validation datasets and kernel density estimation with a Gaussian kernel and bandwidth selected according to Scott’s Rule (Scott1992) on the conditional validation datasets. Specifically, in our results (e.g., Figure 4), we show a $5\times 3$ grid of plots in which the first column contains the partially observed spatial fields and the second and third columns contain the conditional univariate and bivariate kernel density estimates respectively for the selected unobserved locations.
Conditional Mean Fields
The empirical conditional mean fields are computed by averaging the $m=4000$ replicates at each of the unobserved locations in the conditional validation dataset. In our results (e.g., Figures S3 and S6), we display a $5\times 3$ grid of plots that contains the partially observed spatial fields in the first column, the true empirical (Gaussian) or an MCMC-approximated (Brown–Resnick) conditional mean field in the second column, and the empirical NCS conditional mean field in the third column.
S3 Gaussian Process: Tractable Case Study
We applied NCS to a Gaussian process with zero mean and exponential covariance function parametrized by the length scale $\ell>0$ and the variance $\tau^{2}>0$. Hence, the model is
$h$ $\sim\textrm{GP}(0,k(\bm{s}_{1},\bm{s}_{2})),\quad\textrm{where}\quad k(\bm{s}_{1},\bm{s}_{2})=\tau^{2}\exp(-\left.\left\lVert\bm{s}_{1}-\bm{s}_{2}\right\rVert_{2}\middle/\ell\right.),\textrm{ }\bm{s}_{1},\bm{s}_{2}\in\mathcal{S}.$ (S13)
S3.1 Training Details
Parameter U-Net
For the parameter U-Net in the Gaussian process case, we fixed the variance to $\tau^{2}=1.5$ and the proportion of observed locations to $\rho=0.05$, but let the length scale $\ell\in[1.0,5.0]$ vary. To simulate training data, we follow the process described in Section S2.2. Specifically, for training, we simulated $p=1000$ parameters samples from $\tilde{\Theta}=[0.5,5.5]\times\{1.5\}$, $s$ = 1 full spatial field realizations per parameter sample, and $m=100$ masks and timesteps per full spatial field realization.
Proportion U-Net
For the proportion U-Net in the Gaussian process case, we fixed the variance to $\tau^{2}=1.5$, the length scale to $\ell=3$, while we allowed the proportion of observed locations $\rho\in[0.01,0.5]$ to vary. We follow the process in Section S2.2 for simulating training data. Specifically, for training, we sampled $r=50$ observed proportions $\rho$ uniformly from $[0.01,0.525]$, $p=1$ parameters uniformly from $\Theta=\tilde{\Theta}=\{3\}\times\{1.5\}$ for each $\rho$, $s=25$ full spatial field realizations for each $\rho$, and used $m=100$ masks and timesteps for each full spatial field realization.
S3.2 Validation Metrics Particular to the Gaussian Process Case Study
Conditional Correlation Metrics
To assess how well the NCS simulations capture the spatial correlations induced by the true predictive distributions in the Gaussian case study, we produce empirical conditional correlation heatmaps (Figure S4). For a fixed unobserved spatial location in a partially observed spatial field from the conditional validation dataset, we compute the Pearson correlation coefficient (PCC) between the unobserved spatial location and all other unobserved spatial locations using both NCS simulations and conditional simulations from the true predictive distribution. PCC measures the linear correlation between the process evaluated at two unobserved spatial locations $\bm{s}_{1},\bm{s}_{2}$ given the partially observed spatial field. If the PCC is 1-1 or $1$, there is perfect linear correlation given the partially observed spatial field, while a PCC value of 0 indicates no linear correlation given the partially observed field. Specifically, we display a $5\times 2$ grid of plots for both the parameter and proportion U-Nets, in which the first column contains the true empirical correlation heat maps and the second column contains the heatmaps derived from NCS.
S3.3 Results
Visualizations
In Figure S1, we show simulations from the true predictive distributions and NCS. These simulations display similar spatial patterns and seamlessly incorporate the partially observed field as intended. Additionally, amortization appears to be highly effective: as the length scale increases in the left panel, the spatial correlations increase in the NCS simulations, mirroring the true conditional simulations. Further, no matter the number of observed spatial locations, NCS seamlessly integrates the simulated unobserved values with the partially observed field as shown in the right panel. From these results, we conclude that NCS simulations are reasonable as one cannot discriminate between the truth and NCS by the naked eye.
Figure S1: Left panel: NCS with a Gaussian process with variance $\tau^{2}=1.5$, length scale $\ell\in\{1,2,3,4,5\}$ (from top to bottom), and proportion of observed locations $\rho=0.05$. Right panel: NCS with a Gaussian process with variance $\tau^{2}=1.5$, length scale $\ell=3$, and observed proportions $\rho\in\{0.01,0.05,0.1,0.25,0.5\}$ (from top to bottom). For each panel: Left columns: Observations. Middle columns: Underlying spatial fields (generated using unconditional simulation, only partially observed). Right columns: Conditional simulations with NCS implemented with a parameter U-Net (left panel) or a proportion U-Net (right panel).
Conditional Marginal and Bivariate Densities
The true predictive distributions of a Gaussian process are fully determined by its first and second moments. As such, if all the true and NCS conditional univariate and bivariate densities are equivalent up to some small approximation error due to finite sample estimation, then the NCS simulations are representative of the true predictive distributions. In Figure S2, the true and NCS conditional univariate and bivariate densities appear equivalent up to approximation error. While Figure S2 only shows densities for a select number of spatial locations and partially observed spatial fields, we compared conditional univariate and bivariate densities for several other spatial locations and observed similar behavior. Thus, we conclude that the first and second moments of the NCS and true predictive distributions are approximately equivalent, and that the NCS simulations are representative of the true predictive distributions.
Figure S2: Left panel: Predictive densities of a Gaussian process with variance $\tau^{2}=1.5$, length scale $\ell\in\{1,2,3,4,5\}$ (from top to bottom), and observed proportion $\rho=0.05$. Right panel: As in left panel, but with $\tau^{2}=3$, and with observed proportions $\rho\in\{0.01,0.05,0.1,0.25,0.5\}$ (from top to bottom). For each panel: Left columns: Observations and prediction locations for which univariate densities (at the red crosses) and bivariate densities (across the red stars) are visualized. Middle columns: Exact (blue) and NCS (orange) empirical conditional univariate densities at the locations marked with crosses. Right columns: Exact (blue) and NCS (orange) empirical conditional bivariate densities across the locations marked with stars.
Conditional Mean and Correlation Metrics
Figure S3 and Figure S4 contain the true and NCS conditional mean fields and correlation heatmaps, respectively. These serve as additional diagnostics to assess whether—and, if so, how—the first and second moments of the NCS and true predictive distributions differ. In both figures, there are little to no visual differences between the true and NCS empirical conditional mean fields and correlation heatmaps. For both the NCS and true predictive distributions in Figure S4, spatial (conditional) correlations generally increase with the length scale, unless there are observed locations (marked in white) near the spatial location of interest (marked in yellow). This behavior is best illustrated in the fourth row of the left panel; the conditional correlations sharply decrease above the spatial location of interest (in yellow) due to the presence of an observation two pixels above (in white). Yet, the conditional correlations are unaffected below the spatial location of interest due to the absence of observed data in that area. Such behavior is expected for the true predictive distributions; it is yet a striking indication how effective NCS is at correctly conditioning on the given partially observed spatial fields.
Figure S3: Left panel: NCS with a Gaussian process with variance $\tau^{2}=1.5$, length scale $\ell\in\{1,2,3,4,5\}$} (from top to bottom), and observed proportions $\rho=0.05$. Right panel: NCS with a Gaussian process with variance $\tau^{2}=1.5$, length scale $\ell=3$, and observed proportions $\rho\in\{0.01,0.05,0.1,0.25,0.5\}$} (from top to bottom). For each panel: Left columns: Observations. Middle columns: True empirical conditional mean field. Right columns: empirical NCS conditional mean field using a parameter U-Net (left panel) or a proportion U-Net (right panel).
Figure S4: True (left) and NCS (right) empirical conditional correlation heatmaps for location of interest (in yellow) for a Gaussian process with $\tau^{2}=1.5$ and $\ell=1,2,3,4,5$ (top to bottom, left panel) and $\ell=3$ and observed proportions $\rho=.01,.05,.1,.25,.5$ (top to bottom, right panel).
S4 Brown–Resnick: Intractable Case Study
S4.1 Details on Standard Approximation for the Brown–Resnick Process
Further Description
For both FCS and LCS, we utilize the function condrmaxstab in the SpatialExtremes package to approximate the predictive distributions (SpatialExtremes). This function applies a three-step procedure in which the first step involves identifying feasible partitions of the $k$ observed locations based on whether they share common extremal functions (Dombry2012). The extremal functions $\varphi_{\bm{s}_{i}}^{+}$ of a max-stable realization $\bm{z}$ are functions $\eta_{i}W_{i}(\bm{s})$ from the stochastic representation of deHaan1985 in (S9) over the set of spatial locations $\mathcal{S}$ for which the realization $\bm{z}$ and the extremal function $\varphi_{\bm{s}_{i}}^{+}$ evaluated at location $\bm{s}_{i}$ are equal. Extremal functions are not necessarily unique for all spatial locations $\bm{s}_{i}\in\mathcal{S}$ because an extremal function $\varphi_{\bm{s}_{i}}^{+}$ evaluated at another location $\bm{s}_{j}$ may equal the realization $\bm{z}$ at location $\bm{s}_{j}$ as well. As such, this first step requires an approximation via a MCMC algorithm, specifically a Gibbs sampler, to achieve computational tractability because the number of all possible partitions for $k$ locations is the Bell number $B_{k}$ which grows more than exponentially with respect to $k$ (Castruccio). The second and third steps involve simulating extremal and sub-extremal functions respectively from tractable, closed-form distributions and as such, are less computationally intensive. See Dombry2012 for full mathematical details and Huser2019 for another application and assessment of the above Gibbs sampler.
As shown in Figure S5 below, FCS suffers from negative bias for unobserved locations far away from the spatial domain center when the range parameter is small. This is probably due to the insufficient number of subextremal functions drawn by the condrmaxstab function, which affects mostly locations far away from the observed locations when the variogram increases to infinity.
Implementation
To implement FCS and LCS, we use the condrmaxstab function with the given set of observed locations (FCS) or the nearest seven neighbors to the select unobserved location or locations (LCS), the nugget set to $0.00001$, the burn-in period set to 1000, and the thinning factor set to $100$. Here, a nugget of $0.00001$ adds an insignificant amount of noise to the Brown–Resnick unconditional simulations used during the third step, but helps address computational problems which occur in the absence of a nugget.
We use univariate and bivariate LCS to produce conditional univariate and bivariate densities and conditional mean fields. In the univariate case, for LCS, we condition on the seven observations closest to the location of interest. In the bivariate case, we first identify the three nearest neighbors for each of the two locations of interest, and then randomly sample from the remaining unique seven nearest neighbors for each of the two select locations until a conditioning set of size seven is obtained. This process ensures there are neighbors for each of the two locations of interest in the conditioning set for bivariate LCS.
S4.2 Training Details
The major difference between the Gaussian and Brown–Resnick case studies is the log-transformation of the training data simulated from the Brown–Resnick process. As such, NCS produces conditional simulations on the (log-transformed) Gumbel scale, and all results are on the Gumbel scale.
Parameter U-Net
The details for simulating training are the same as for the Parameter U-Net in the Gaussian process case study (Section S3.1) except that the number of proportion values was set to $r=200$ for the training data.
Proportion U-Net
The details for simulating training data are the same as for the Proportion U-Net in the Gaussian process case study.
Small Conditioning Set U-Net
@@
We train U-Nets for the small conditioning sets via the following data generation process: We first fix the parameter $\bm{\theta}=(\lambda,\nu)^{\prime}$ and select a minimum $o_{L}=1$ and maximum $o_{U}=10$ number of observations for the conditioning sets according to the experimental settings. For each data draw, we iterate through each number of observed locations $o_{j}$ starting from the minimum $o_{L}$ and ending with the maximum $o_{U}$. For each observed number $o_{j}$, we simulate $s$ full spatial field realizations via the unconditional simulator with fixed parameter $\bm{\theta}$. For each of these $s$ full spatial fields, we construct $m$ masks and simulate $m$ discrete timesteps uniformly from $\{0,1,\dots,T\}$. Each mask is constructed from a conditioning set $\mathring{\mathcal{S}}$ containing $o_{j}$ locations uniformly sampled without replacement from $\mathcal{S}$. Algorithm S2 illustrates this data process, and the resulting data have the form shown in (S12).
Algorithm S2 Training and Validation Data for Small Conditioning Set U-Net